# Unbounded-width CSPs are Untestable in a Sublinear Number of Queries

### v1

@@ 1 Introduction
Given a CSP template $(D,\Gamma)$, we define $\mathrm{CSP}(\Gamma)$ as the (infinite) collection of all instances $I=(V,\mathcal{C})$, where $V$ is a finite set of variables and $\mathcal{C}=\{C_{1},\dots,C_{m}\}$ is a multi-set of constraints satisfying the following. Each constraint $C_{i}$ is a pair $(\mathbf{v},R)$, where $\mathbf{v}=(v_{1},\dots,v_{k})$ is a tuple of distinct [^1] Some authors allow repetition of variables in a constraint. In our setting, one can also “artificially” allow repetitions by adding more relations to the relational structure. variables from $V$, and $R:D^{k}\to\{0,1\}$ is a relation from $\Gamma$. The variable set $\{v_{1},\dots,v_{k}\}\subseteq V$ is called the scope of $C_{i}$.
@@ 1.1 Testing CSP Satisfiability
The bounded-degree query model was first introduced by [GR02], in the context of graph property testing. One of the earliest results in this model is that testing bipartiteness (i.e. testing satisfiability of $\mathsf{2COL}$ instances) requires at least $\Omega(\sqrt{n})$ queries [GR02] (for some degree bound $d$ and error $\varepsilon>0$) and can be done with at most $\widetilde{O}(\sqrt{n})$ queries [GR99] (for any degree bound $d$ and any error $\varepsilon>0$). In contrast, Bogdanov, Obata, and Trevisan [BOT02] showed that testing vertex 3-colorability requires $\Omega(n)$ queries. [^2] These seminal works, along with a large number of subsequent studies (such as [GR11, BSS10, YI10, YI12, CMOS19], to name a few), also investigated a wide range of other graph properties within the bounded-degree model, many of which are not constraint satisfaction problems.
The study of testing CSP satisfiability in the bounded-degree model was already formally initiated in the early work of Bogdanov, Obata, and Trevisan [BOT02], who showed that, beyond the specific case of $\mathsf{3COL}$, several other CSPs (such as $\mathsf{3SAT}$ and $\mathsf{3LIN}$) require $\Omega(n)$ queries to test in this model. A more systematic investigation was later carried out by Yoshida [Yos11], who proved that any CSP robustly solved by BasicLP (the basic linear program) has a satisfiability tester that makes a constant number of queries. Typical examples of such CSPs include $\mathsf{1SAT}$ and $\mathsf{3HornSAT}$. For the satisfiability testing of any CSP that is not solved by BasicLP[Yos11] further showed that at least $\Omega(\sqrt{n})$ queries are necessary. Examples of such CSPs include $\mathsf{2SAT}$ and $\mathsf{2COL}$. [^3] Robust solvability by BasicLP is a stronger condition than solvability by BasicLP, so [Yos11] left open the possibility of CSPs with superconstant but $o(\sqrt{n})$ query complexity. However, a recent unpublished note by Brady [Bra22] proves that all CSPs solvable by BasicLP are also robustly solvable by it, thereby turning the result of [Yos11] into a full dichotomy theorem.
@@
The notion of bounded width (formally defined in Appendix A) originates from [FV98] and has been extensively studied in the universal-algebraic approach to CSPs. Bounded-width CSPs can be solved in polynomial time (see Proposition A.2), so the class of unbounded-width CSPs includes all CSPs whose satisfiability is NP-hard to decide (unless $\textbf{P}=\textbf{NP}$, in which case all CSPs are trivially NP-hard). Consequently, Theorem 1.6 implies that all CSPs known to be NP-hard are also unconditionally (and maximally) hard to test in the bounded-degree model. This encompasses, for example, the property of $k$-colorability in $\ell$-uniform hypergraphs for all $k,\ell\geqslant 2$ with $(k,\ell)\neq(2,2)$ (see Remark A.5). Our result thus unifies and extends several prior lower bounds from works such as [BOT02, YI10, ACC25]. [^4] The recent work of [ACC25] proves that testing $k$-colorability in $k$-uniform hypergraphs requires $\Omega(n)$ queries, for any $k\geqslant 3$.
@@
A property tester can be viewed as an algorithm that approximates the distance of an input to satisfying a property — albeit in a very coarse sense, as it only distinguishes zero distance from sufficiently large distance. Nevertheless, many property testers (for example, the constant-query tester of [Yos11] for certain CSPs) operate by explicitly estimating this distance and checking whether it is close to zero. On the negative side, a hardness result for property testing (such as our Theorem 1.6) can often be interpreted as a hardness of approximation result. This connection to approximation algorithms was, in fact, one of the original motivations in [GGR98], where graph property testing was first introduced.
@@ 1.2.1 MaxCSP Problems
The complexity of MaxCSP problems has been extensively studied in the context of polynomial-time algorithms, and a central question is to determine for which tuples $(\Gamma,c,s)$ the problem $\mathsf{MaxCSP}(\Gamma)[c,s]$ can be solved in polynomial time. Notably, Raghavendra [Rag08] almost [^6] Except for the perfect-completeness case (i.e. the case $c=1$), which remains a mystery. characterizes all such tuples, conditioned on the Unique Games Conjecture of [Kho02].
@@
We argue that exactly due to this difference in instance regimes, the bounded-degree model provides a stronger connection to the study of polynomial-time algorithms for MaxCSPs, at least in the respect of worst-case approximation ratios. First, MaxCSP problems on general instances can often be reduced to bounded-degree instances (see, e.g., [Tre01]), so bounded-degree instances already capture much of the hardness in the problems. Second, MaxCSP problems tend to be easier on dense instances (see, e.g., [AKK95]). Indeed, [AS03] shows that every CSP admits a constant-query satisfiability tester in the dense graph model. In contrast, in the bounded-degree model, satisfiability testing already exhibits a wide spectrum of complexities across CSP templates, with query complexity ranging from constant, to $\widetilde{\Theta}(\sqrt{n})$, up to $\Theta(n)$.
@@
In the context of polynomial-time algorithms, for certain CSP templates $(D,\Gamma)$, the perfect-completeness problem $\mathsf{MaxCSP}(\Gamma)[1,1-\varepsilon]$ can exhibit a markedly different level of complexity from the (potentially harder) imperfect-completeness problem $\mathsf{MaxCSP}(\Gamma)[1-\varepsilon^{\prime},1-\varepsilon]$, even when $\varepsilon^{\prime}\to 0$ and $\varepsilon$ is fixed. For example, in the case of $\mathsf{3LIN}$ (an unbounded-width CSP), the former problem admits a polynomial-time algorithm via Gaussian elimination, whereas the latter is $\mathbf{NP}$-hard (due to [Hås01]). More generally, Dalmau and Krokhin [DK13] proved that the imperfect-completeness problem is $\mathbf{NP}$-hard for all unbounded-width CSP templates.
Theorem 1.9 ([DK13]).
@@
In terms of techniques, however, the proof of Theorem 1.8 runs in direct parallel to that of Theorem 1.9 in [DK13]. While [DK13] constructs a reduction from the $\mathbf{NP}$-hardness of approximating $\mathsf{3LIN}$ instances (established by [Hås01]), our proof of Theorem 1.8 builds an analogous reduction from the linear-query lower bound for $\mathsf{3LIN}$ instances (proved in [BOT02]). Unlike in the $\mathbf{NP}$ setting, where a reduction need only run in polynomial time, our bounded-degree query setting requires addressing several additional subtleties. In particular, certain parts of the proof (for example, Section 5.1) rely on a careful combination of techniques from both [BOT02] and [DK13].
@@
The result of [DK13] (Theorem 1.9) was later shown to be tight by Barto and Kozik [BK16]. Specifically, they proved that for every CSP template $(D,\Gamma)$ of bounded width and every constant $\varepsilon\in(0,1)$, there exists a constant $\varepsilon^{\prime}\in(0,\varepsilon)$ such that the problem $\mathsf{MaxCSP}(\Gamma)[1-\varepsilon^{\prime},1-\varepsilon]$ belongs to $\mathbf{P}$. Together with the result of [DK13], this establishes that bounded width exactly characterizes the class of robustly solvable CSPs, as conjectured earlier by Guruswami and Zhou [GZ12].
@@
The past decades have seen a growing body of work on MaxCSP problems in sublinear-space algorithmic models (such as [KKS14, KK19, CGSV24, CGS+22, SSSV25, FMW25a], to name a few), particularly in the contexts of sketching and streaming algorithms. Recent studies [FMW25b, FMW25a] reveal a phenomenological connection between the multi-pass streaming model and the bounded-degree query model: the class of MaxCSP problems that are easy in the bounded-degree model (admitting algorithms with constantly many queries) almost [^7] Except for the perfect-completeness case, which remains a mystery for both models. coincides with those that are easy in the multi-pass streaming model (admitting algorithms with constantly many passes and polylogarithmic space). We view this connection as further motivation for studying MaxCSP and other approximation problems in the bounded-degree query model.
@@
We remark that, as discussed in [FMW25a], query lower bounds in sublinear time models are usually easier to establish than space lower bounds, and thus Theorem 1.8 can be seen as a preliminary step towards answering 1.11.
@@ 1.3 Open Problems
In their original work, Bogdanov, Obata and Trevisan [BOT02] asked whether, given a CSP template $(D,\Gamma)$, one can determine the optimal hardness gap $\varepsilon$ for which $\mathsf{MaxCSP}(\Gamma)[1,1-\varepsilon]$ requires $\Omega(n)$ queries. This question remains wide open (even for “simple” templates such as $\mathsf{3COL}$). Here, we pose the following more ambitious version.
@@ 2 Preliminaries
As mentioned in Section 1.2.3, the proof of Theorem 1.8 is by constructing a reduction from “linear equation” instances. As is the case in [BOT02], the reduction is centered around the fact that the linear equations can be “simulated” by some constant-size gadgets built from constraints in the target template $(D,\Gamma)$. The following general definition captures such gadget constructions.
@@
Over the past decades, the development of the powerful framework of universal algebra has enabled a systematic study of the power and limitations of gadget constructions for CSPs. In contrast to the case-by-case constructions used in earlier works such as [BOT02, YI10, ACC25], we draw on established results in universal algebra to fully exploit the strength of gadget constructions. The universal-algebraic result (Theorem 3.2) used in the proof of Theorem 1.8 is highly nontrivial (though by now perhaps standard) and serves as the main workhorse that allows us to unify the existing $\Omega(n)$ query lower bounds. The remainder of this section is devoted to introducing the necessary background in universal algebra before formally stating the lemma in the next section.
@@ 2.1 Relational Structures
We begin by introducing some basic terminology for studying relational structures, as defined in Definition 1.1, with an emphasis on maps between different relational structures.
@@
Every relational structure $(D,\Gamma)$ is homomorphically equivalent to a core.
@@
We claim that $(D_{1},\Gamma_{1})$ is a core. Otherwise, there exists a relational endomorphism $\varphi_{1}$ on $(D_{1},\Gamma_{1})$ that is not bijective. In particular, its image is a proper subset $D_{2}\subsetneq D_{1}$. But then the composition $\varphi_{1}\circ\varphi$ defines a homomorphism from $D$ to $D_{2}$, contradicting the minimality of $D_{1}$. Hence $(D_{1},\Gamma_{1})$ must be a core. ∎

### v2

@@ 1 Introduction
Given a CSP template $(D,\Gamma)$, we define $\mathrm{CSP}(\Gamma)$ as the (infinite) collection of all instances $I=(V,\mathcal{C})$, where $V$ is a finite set of variables and $\mathcal{C}=\{C_{1},\dots,C_{m}\}$ is a multi-set of constraints satisfying the following. Each constraint $C_{i}$ is a pair $(\mathbf{v},R)$, where $\mathbf{v}=(v_{1},\dots,v_{k})$ is a tuple of (not necessarily distinct) [^1] Some authors allow repetition of variables in a constraint. In our setting, one can also “artificially” allow repetitions by adding more relations to the relational structure. See Section 2.1 for details. variables from $V$, and $R:D^{k}\to\{0,1\}$ is a relation from $\Gamma$. The variable set $\{v_{1},\dots,v_{k}\}\subseteq V$ is called the scope of $C_{i}$.
@@ 1.1 Testing CSP Satisfiability
The bounded-degree query model was first introduced by [goldreich2002property], in the context of graph property testing. One of the earliest results in this model is that testing bipartiteness (i.e. testing satisfiability of $\mathsf{2COL}$ instances) requires at least $\Omega(\sqrt{n})$ queries [goldreich2002property] (for some degree bound $d$ and error $\varepsilon>0$) and can be done with at most $\widetilde{O}(\sqrt{n})$ queries [goldreich1999sublinear] (for any degree bound $d$ and any error $\varepsilon>0$). In contrast, Bogdanov, Obata, and Trevisan [bogdanov2002lower] showed that testing vertex 3-colorability requires $\Omega(n)$ queries. [^2] These seminal works, along with a large number of subsequent studies (such as [goldreich2011testing, benjaminia2010every, yoshida2010query, yoshida2012property, czumaj2019planar], to name a few), also investigated a wide range of other graph properties within the bounded-degree model, many of which are not constraint satisfaction problems.
The study of testing CSP satisfiability in the bounded-degree model was already formally initiated in the early work of Bogdanov, Obata, and Trevisan [bogdanov2002lower], who showed that, beyond the specific case of $\mathsf{3COL}$, several other CSPs (such as $\mathsf{3SAT}$ and $\mathsf{3LIN}$) require $\Omega(n)$ queries to test in this model. A more systematic investigation was later carried out by Yoshida [yoshida2011optimal], who proved that any CSP robustly solved by BasicLP (the basic linear program) has a satisfiability tester that makes a constant number of queries. Typical examples of such CSPs include $\mathsf{1SAT}$ and $\mathsf{3HornSAT}$. For the satisfiability testing of any CSP that is not solved by BasicLP[yoshida2011optimal] further showed that at least $\Omega(\sqrt{n})$ queries are necessary. Examples of such CSPs include $\mathsf{2SAT}$ and $\mathsf{2COL}$. [^3] Robust solvability by BasicLP is a stronger condition than solvability by BasicLP, so [yoshida2011optimal] left open the possibility of CSPs with superconstant but $o(\sqrt{n})$ query complexity. However, a recent unpublished note by Brady [brady2022notes] proves that all CSPs solvable by BasicLP are also robustly solvable by it, thereby turning the result of [yoshida2011optimal] into a full dichotomy theorem.
@@
The notion of bounded width (formally defined in Appendix A) originates from [feder1998computational] and has been extensively studied in the universal-algebraic approach to CSPs. Bounded-width CSPs can be solved in polynomial time (see Proposition A.2), so the class of unbounded-width CSPs includes all CSPs whose satisfiability is NP-hard to decide (unless $\textbf{P}=\textbf{NP}$, in which case all CSPs are trivially NP-hard). Consequently, Theorem 1.6 implies that all CSPs known to be NP-hard are also unconditionally (and maximally) hard to test in the bounded-degree model. This encompasses, for example, the property of $k$-colorability in $\ell$-uniform hypergraphs for all $k,\ell\geqslant 2$ with $(k,\ell)\neq(2,2)$ (see Remark A.5). Our result thus unifies and extends several prior lower bounds from works such as [bogdanov2002lower, yoshida2010query, aaronson2025property]. [^4] The recent work of [aaronson2025property] proves that testing $k$-colorability in $k$-uniform hypergraphs requires $\Omega(n)$ queries, for any $k\geqslant 3$.
@@
A property tester can be viewed as an algorithm that approximates the distance of an input to satisfying a property — albeit in a very coarse sense, as it only distinguishes zero distance from sufficiently large distance. Nevertheless, many property testers (for example, the constant-query tester of [yoshida2011optimal] for certain CSPs) operate by explicitly estimating this distance and checking whether it is close to zero. On the negative side, a hardness result for property testing (such as our Theorem 1.6) can often be interpreted as a hardness of approximation result. This connection to approximation algorithms was, in fact, one of the original motivations in [goldreich1998property], where graph property testing was first introduced.
@@ 1.2.1 MaxCSP Problems
The complexity of MaxCSP problems has been extensively studied in the context of polynomial-time algorithms, and a central question is to determine for which tuples $(\Gamma,c,s)$ the problem $\mathsf{MaxCSP}(\Gamma)[c,s]$ can be solved in polynomial time. Notably, Raghavendra [raghavendra2008optimal] almost [^6] Except for the perfect-completeness case (i.e. the case $c=1$), which remains a mystery. characterizes all such tuples, conditioned on the Unique Games Conjecture of [khot2002power].
@@
We argue that exactly due to this difference in instance regimes, the bounded-degree model provides a stronger connection to the study of polynomial-time algorithms for MaxCSPs, at least in the respect of worst-case approximation ratios. First, MaxCSP problems on general instances can often be reduced to bounded-degree instances (see, e.g., [trevisan2001non]), so bounded-degree instances already capture much of the hardness in the problems. Second, MaxCSP problems tend to be easier on dense instances (see, e.g., [arora1995polynomial]). Indeed, [alon2003testing] shows that every CSP admits a constant-query satisfiability tester in the dense graph model. In contrast, in the bounded-degree model, satisfiability testing already exhibits a wide spectrum of complexities across CSP templates, with query complexity ranging from constant, to $\widetilde{\Theta}(\sqrt{n})$, up to $\Theta(n)$.
@@
In the context of polynomial-time algorithms, for certain CSP templates $(D,\Gamma)$, the perfect-completeness problem $\mathsf{MaxCSP}(\Gamma)[1,1-\varepsilon]$ can exhibit a markedly different level of complexity from the (potentially harder) imperfect-completeness problem $\mathsf{MaxCSP}(\Gamma)[1-\varepsilon^{\prime},1-\varepsilon]$, even when $\varepsilon^{\prime}\to 0$ and $\varepsilon$ is fixed. For example, in the case of $\mathsf{3LIN}$ (an unbounded-width CSP), the former problem admits a polynomial-time algorithm via Gaussian elimination, whereas the latter is $\mathbf{NP}$-hard (due to [haastad2001some]). More generally, Dalmau and Krokhin [dalmau2013robust] proved that the imperfect-completeness problem is $\mathbf{NP}$-hard for all unbounded-width CSP templates.
Theorem 1.9 ([dalmau2013robust]).
@@
In terms of techniques, however, the proof of Theorem 1.8 runs in direct parallel to that of Theorem 1.9 in [dalmau2013robust]. While [dalmau2013robust] constructs a reduction from the $\mathbf{NP}$-hardness of approximating $\mathsf{3LIN}$ instances (established by [haastad2001some]), our proof of Theorem 1.8 builds an analogous reduction from the linear-query lower bound for $\mathsf{3LIN}$ instances (proved in [bogdanov2002lower]). Unlike in the $\mathbf{NP}$ setting, where a reduction need only run in polynomial time, our bounded-degree query setting requires addressing several additional subtleties. In particular, certain parts of the proof (for example, Section 5.1) rely on a careful combination of techniques from both [bogdanov2002lower] and [dalmau2013robust].
@@
The result of [dalmau2013robust] (Theorem 1.9) was later shown to be tight by Barto and Kozik [barto2016robustly]. Specifically, they proved that for every CSP template $(D,\Gamma)$ of bounded width and every constant $\varepsilon\in(0,1)$, there exists a constant $\varepsilon^{\prime}\in(0,\varepsilon)$ such that the problem $\mathsf{MaxCSP}(\Gamma)[1-\varepsilon^{\prime},1-\varepsilon]$ belongs to $\mathbf{P}$. Together with the result of [dalmau2013robust], this establishes that bounded width exactly characterizes the class of robustly solvable CSPs, as conjectured earlier by Guruswami and Zhou [guruswami2012tight].
@@
The past decades have seen a growing body of work on MaxCSP problems in sublinear-space algorithmic models (such as [kapralov2014streaming, kapralov2019optimal, chou2024sketching, chou2022linear, saxena2025streaming, fei2025dichotomy], to name a few), particularly in the contexts of sketching and streaming algorithms. Recent studies [fei2025multi, fei2025dichotomy] reveal a phenomenological connection between the multi-pass streaming model and the bounded-degree query model: the class of MaxCSP problems that are easy in the bounded-degree model (admitting algorithms with constantly many queries) almost [^7] Except for the perfect-completeness case, which remains a mystery for both models. coincides with those that are easy in the multi-pass streaming model (admitting algorithms with constantly many passes and polylogarithmic space). We view this connection as further motivation for studying MaxCSP and other approximation problems in the bounded-degree query model.
@@
We remark that, as discussed in [fei2025dichotomy], query lower bounds in sublinear time models are usually easier to establish than space lower bounds, and thus Theorem 1.8 can be seen as a preliminary step towards answering 1.11.
@@ 1.3 Open Problems
In their original work, Bogdanov, Obata and Trevisan [bogdanov2002lower] asked whether, given a CSP template $(D,\Gamma)$, one can determine the optimal hardness gap $\varepsilon$ for which $\mathsf{MaxCSP}(\Gamma)[1,1-\varepsilon]$ requires $\Omega(n)$ queries. This question remains wide open (even for “simple” templates such as $\mathsf{3COL}$). Here, we pose the following more ambitious version.
@@ 2 Preliminaries
As mentioned in Section 1.2.3, the proof of Theorem 1.8 is by constructing a reduction from “linear equation” instances. As is the case in [bogdanov2002lower], the reduction is centered around the fact that the linear equations can be “simulated” by some constant-size gadgets built from constraints in the target template $(D,\Gamma)$. The following general definition captures such gadget constructions.
@@
Over the past decades, the development of the powerful framework of universal algebra has enabled a systematic study of the power and limitations of gadget constructions for CSPs. In contrast to the case-by-case constructions used in earlier works such as [bogdanov2002lower, yoshida2010query, aaronson2025property], we draw on established results in universal algebra to fully exploit the strength of gadget constructions. The universal-algebraic result (Theorem 3.2) used in the proof of Theorem 1.8 is highly nontrivial (though by now perhaps standard) and serves as the main workhorse that allows us to unify the existing $\Omega(n)$ query lower bounds. The remainder of this section is devoted to introducing the necessary background in universal algebra before formally stating the lemma in the next section.
@@ 2.1 Relational Structures
We begin by introducing some basic terminology for studying relational structures.
Recall from Definition 1.2 that we do not allow repetition of variables in individual constraints of CSP instances. For example, if E3Sat is the collection of relations $R:\{0,1\}^{3}\rightarrow\{0,1\}$ such that $|R^{-1}(1)|=7$, then $\mathrm{CSP}(\textsc{E3Sat})$ is the collection of all CNF formulas where each clause contains exactly 3 literals. In order to take into account CNF formulas with some clauses containing less than 3 literals (as is standard in the $\mathsf{3SAT}$ problem), we enlarge the relation set E3Sat into 3Sat — the collection of relations $R:\{0,1\}^{k}\rightarrow\{0,1\}$ such that $|R^{-1}(1)|=2^{k}-1$ and $k\in\{1,2,3\}$. Now $\mathrm{CSP}(\textsc{3Sat})$ expresses the standard version of the $\mathsf{3SAT}$ problem.
Given a general relation set $\Gamma$, in order to allow variable repetition in $\mathrm{CSP}(\Gamma)$ instances, we can perform the following operation on $\Gamma$.
Definition 2.3.
For any relation $R:D^{k}\rightarrow\{0,1\}$ on a finite domain $D$, we let $\overline{\{R\}}$ be the collection of relations $R^{\prime}:D^{\ell}\rightarrow\{0,1\}$ such that there exists a surjective map $\pi:[k]\rightarrow[\ell]$ and $R^{\prime}(x_{1},\dots,x_{\ell})=R(x_{\pi(1)},\dots,x_{\pi(k)})$ for all $x_{1},\dots,x_{\ell}\in D$. For any relational structure $(D,\Gamma)$, we let $\overline{\Gamma}$ be the union $\bigcup_{R\in\Gamma}\overline{\{R\}}$. A relational structure $(D,\Gamma)$ is said to be repetition-closed [^8] This is not standard terminology. if $\Gamma=\overline{\Gamma}$.
By a standard reduction, in our main theorem (Theorem 1.8) it suffices to consider relational structures that are repetition-closed. Formally, it will be proved in Appendix B that the following statement implies Theorem 1.8.
Theorem 2.4.
For any repetition-closed CSP template $(D,\Gamma)$ of unbounded width, there exist constants $\varepsilon,\alpha\in(0,1)$ and $d\in\mathbb{N}$ such that any algorithm for $\mathsf{MaxCSP}(\Gamma)[1,1-\varepsilon]$ in the $\mathrm{BD}^{*}(d,\alpha,n)$ model must make $\Omega(n)$ queries.
We next turn to study maps between relational structures.
@@
Every relational structure $(D,\Gamma)$ is homomorphically equivalent to a core. If $(D,\Gamma)$ is repetition-closed, then $(D,\Gamma)$ is homomorphically equivalent to a repetition-closed core.
@@
We claim that $(D_{1},\Gamma_{1})$ is a core. Otherwise, there exists a relational endomorphism $\varphi_{1}$ on $(D_{1},\Gamma_{1})$ that is not bijective. In particular, its image is a proper subset $D_{2}\subsetneq D_{1}$. But then the composition $\varphi_{1}\circ\varphi$ defines a homomorphism from $D$ to $D_{2}$, contradicting the minimality of $D_{1}$. Hence $(D_{1},\Gamma_{1})$ must be a core.
Now assume $(D,\Gamma)$ is repetition closed. Any relation in $\Gamma_{1}$ of arity $k$ is the restriction of some relation $R:D^{k}\rightarrow\{0,1\}$ to $D_{1}$. Therefore, for any surjection $\pi:[k]\rightarrow[\ell]$, there exists $R^{\prime}\in\Gamma$ such that $R^{\prime}(x_{1},\dots,x_{\ell})=R(x_{\pi(1)},\dots,x_{\pi(k)})$ for all $x_{1},\dots,x_{\ell}\in D$, and in particular $R^{\prime}_{|D_{1}}(x_{1},\dots,x_{\ell})=R_{|D_{1}}(x_{\pi(1)},\dots,x_{\pi(k)})$ for all $x_{1},\dots,x_{\ell}\in D_{1}$. As $R^{\prime}_{|D_{1}}\in\Gamma_{1}$, we conclude that $(D_{1},\Gamma_{1})$ is also repetition closed. ∎